Parth (parthkohli):

I'm asked to prove something: If \(\sin\theta + \cos\theta = \sqrt{2}\sin\theta\) then show that \(\cos\theta - \sin\theta = \sqrt{2}\cos\theta\). I don't even think that's correct, because multiplying both equations, I get \(\cos^2\theta - \sin^2\theta = 2\sin\theta\cos\theta \Rightarrow \cos(2\theta) = \sin(2\theta)\). That's wrong. Totally wrong.

Parth (parthkohli):

But there are cases where \(\cos(2\theta) = \sin(2\theta)\). Well...

Parth (parthkohli):

@ajprincess Please help :P

Parth (parthkohli):

I'm thinking of squaring both the sides.

OpenStudy (whpalmer4):

None of the first 4 solutions where \(\cos(2\theta) = \sin(2\theta)\) appear to satisfy the equation you are asked to prove...

Parth (parthkohli):

Yay, thanks! :-)